# Trajeglish: Learning the Language of Driving Scenarios

### 07 Dec 2023 v1

@@ Abstract
A longstanding challenge for self-driving development is simulating dynamic driving scenarios seeded from recorded driving logs. In pursuit of this functionality, we apply tools from discrete sequence modeling to model how vehicles, pedestrians and cyclists interact in driving scenarios. Using a simple data-driven tokenization scheme, we discretize trajectories to centimeter-level resolution using a small vocabulary. We then model the multi-agent sequence of motion tokens with a GPT-like encoder-decoder that is autoregressive in time and takes into account intra-timestep interaction between agents. Scenarios sampled from our model exhibit state-of-the-art realism; our model tops the Waymo Sim Agents Benchmark, surpassing prior work along the realism meta metric by 3.3% and along the interaction metric by 9.9%. We ablate our modeling choices in full autonomy and partial autonomy settings, and show that the representations learned by our model can quickly be adapted to improve performance on nuScenes. We additionally evaluate the scalability of our model with respect to parameter count and dataset size, and use density estimates from our model to quantify the saliency of context length and intra-timestep interaction for the traffic modeling task.
@@ 1 Introduction
In this work, we propose an autoregressive model of the motion of roadusers that can be used to simulate how humans might react if a self-driving system were to choose a given sequence of actions. At test time, as visualized in Fig. 1, the model functions as a policy, outputting a categorical distribution over the set of possible states an agent might move to at each timestep. Iteratively sampling actions from the model results in diverse, scene-consistent multi-agent rollouts of arbitrary length. We call our approach “Trajeglish” due to the fact that we model multi-agent trajectories as a sequence of discrete tokens, similar to the representation used in language modeling, and to make an analogy between how roadusers use vehicle motion to communicate and how people use verbal languages, like English, to communicate.
@@ 1.1 Related Work
Figure 2: Trajeglish Visualizations of samples from our model. Rollouts within each row are given the same single-timestep initialization, outlined in black. Future trajectories become lighter for timesteps farther into the future. While some tracks overlap in the figure, they do not overlap when time is taken into account; there are no collisions in these rollouts. Videos are available on our project page.
@@ 3.1 Tokenization
We propose an easily parallelizable approach for finding template sets with low discretization error. We collect a large number of state transitions observed in data, sample one of them, filter transitions that are within $\epsilon$ meters, and repeat $|V|$ times. Pseudocode for this algorithm is included in Alg. 1. We call this method for sampling candidate templates “k-disks” given it’s similarity to k-means++, the standard algorithm for seeding the anchors k-means (Arthur & Vassilvitskii, 2007), as well as the Poisson disk sampling algorithm (Cook, 1986). We visualize the template sets found using k-disks with minimum discretization error in Fig. 4. We verify in Fig. 5 that the tokenized action distribution is similar on WOMD train and validation despite the fact that the templates are optimized on the training set. We show in Fig. 6 that the discretization error induced by templates sampled with k-disks is in general much better than that of k-means, across agent types. A comprehensive evaluation of k-disks in comparison to baselines is in Sec. A.3.
Figure 7: Trajeglish modeling We train an encoder-decoder transformer that predicts the action token of an agent conditional on previous action tokens, map information, and agent information available at $t=0$. The diagram represents the forward pass of the network during training in which $t=0$ agent information, map objects, and motion tokens are passed into the network using a causal mask and the model is trained to predict the next motion token, shown in the top right.
@@ 3.2 Modeling
Our model takes as input two modalities that encode the initial scene. The first is the initial state of the agents in the scene which includes the length, width, initial position, initial heading, and object class. We apply a single layer MLP to encode these values per-agent to an embedding of size $C$. We then add a positional embedding that encodes the agent’s order as well as agent identity across the action sequence. The second modality is the map. We use the WOMD representation of a map as a collection of “map objects”, where a map object might be a variable-length polyline representing a lane, a sidewalk, a crosswalk, etc.. We apply a VectorNet encoder to encode the map to a sequence of embeddings for at most $M$ map objects (Gao et al., 2020). Note that although the model is not permutation equivariant to the agents, it is permutation invariant to the ordering of the map objects. Similar to Wayformer (Nayakanti et al., 2022), we then apply a layer of latent query attention that outputs a final encoding of the scene initialization.
@@ 4.1 WOMD Sim Agents Benchmark
Figure 8: Partial control ADE Left shows the ADE for the vehicles selected for evaluation under partial control, but for rollouts where the agents are fully autonomous. Right shows the ADE for the same vehicles but with all other agents on replay. When agents controlled by Trajeglish go first in the permutation order, they behave similarly to the no intra model. When they go last, utilize the intra-timestep information to produce interaction more similar to recorded logs, achieving a lower ADE.
@@ 4.2 Ablation
To simplify our ablation study, we test models in this section on the scenarios they train on, of at most 24 agents and 6.4 seconds in length. We compare performance across 5 variants of our model. Both “trajeglish” and “trajeglish w/ reg.” refer to our model, the latter using the noisy tokenization strategy discussed in Sec. A.2. The “no intra” model is an important baseline designed to mimic the behavior of behavior cloning baselines used in Symphony (Igl et al., 2022) and “Imitation Is Not Enough” (Lu et al., 2023). For this baseline, we keep the same architecture but adjust the masking strategy in the decoder to not attend to actions already chosen for the current timestep. The “marginal” baseline is designed to mimic the behavior of models such as Wayformer (Nayakanti et al., 2022) and MultiPath++ (Varadarajan et al., 2021) that are trained to model the distribution over single agent trajectories instead of multi-agent scene-consistent trajectories. For this baseline, we keep the same architecture but apply a mask to the decoder that enforces that the model can only attend to previous actions chosen by the current agent. Our final baseline is the same as the marginal baseline but without a map encoder. We use this baseline to understand the extent to which the models rely on the map for traffic modeling.
@@
We report results in Fig. 8 in a partial controllability setting in which a single agent in each scenario is chosen to be controlled by the traffic model and all other agents are set to replay. The single-agent ADE (average distance error) for the controlled-agent is similar in full autonomy rollouts for all models other than the model that does not condition on the map, as expected. However, in rollouts where all other agents are placed on replay, the replay trajectories leak information about the trajectory that the controlled-agent took in the data, and as a result, the no-intra and trajeglish rollouts have a lower ADE. Additionally, the trajeglish rollouts in which the controlled-agent is placed first do not condition on intra-timestep information and therefore behave identically to the no-intra baseline, whereas rollouts where the controlled-agent is placed last in the order provide the model with more information about the replay trajectories and result in a decreased ADE.
@@ 4.3 Analysis
We visualize the embeddings that the model learns in Fig. 11. Through the task of predicting the next token, the model learns a similarity matrix across tokens that reflects the euclidean distance between the actions the tokens represent.
@@
Figure 11: Token Embedding Visualization We run PCA on the model embeddings at initialization and at convergence, and plot the $(x,y)$ location of each of the token templates using the top 3 principal component values to determine the hue, lightness, and saturation of the point. The model learns that tokens that correspond to actions close together in euclidean space represent semantically similar actions. Note that the heading of each action is not visualized, which also affects action similarity. Additionally, the top 3 principal components include only 35% of the variance, explaining why some colors repeat.
@@ A.3 Tokenization Analysis
$(x,y,h)$-grid - We independently discretize change in longitudinal and latitudinal position and change in heading, and treat the template set as the product of these three sets. For vocabulary sizes of 128/256/384/512 respectively, we use 6/7/8/9 values for $x$ and $y$, and 4/6/7/8 values for $h$. These values are spaced evenly between (-0.3, 3.5) m for $x$, (-0.2 m, 0.2 m) for $y$, and (-0.1, 0.1) rad for $h$.
$(x,y)$-grid - We independenly discretize change in only the location. We choose the heading for each template based on the heading of the state-to-state transition found in the data with a change in location closest to the template location. Compared to the $(x,y,h)$-grid baseline, this approach assumes heading is deterministic given location in order to gain resolution in location. We use 12/16/20/23 values for $x$ and $y$ with the same bounds as in the $(x,y,h)$-grid baseline.
@@
We offer several comparisons between these methods. In Fig. 12, we plot the expected corner distance between trajectories and tokenized trajectories as a function of trajectory length for the template sets found with k-disks. In Fig. 13, we compare the tokenization error as a function of trajectory length and find that grid-based tokenizers create large oscillations. To calibrate to a metric more relevant to the traffic modeling task, we compare the collision rate between raw trajectories as a function of trajectory length for the raw scenarios and the tokenized scenarios using k-disk template sets of size 128, 256, 384, and 512 in Fig. 14. We observe that a vocabulary size of 384 is sufficient to avoid creating extraneous collisions. Finally, Fig. 15 plots the full distribution of discretizaion errors for each of the baselines and Tab. 2 reports the expected discretization error across vocabulary sizes for each of the methods.
@@ A.5 Extended Rollouts for WOMD Sim Agents Benchmark
Table 3: WOMD sim agents validation
@@
A.6 Additional Ablation Results
@@ A.7 Additional Analysis
Context length refers to the number of tokens that the model has to condition on when predicting the distribution over the next token. Intuitively, as the model is given more context, the model should get strictly better at predicting the next token. We quantify this effect in Fig. A.7. We find that the relative decrease in cross entropy from increasing the context length drops off steeply for our model for pedestrians and cyclists, which aligns with the standard intuition that these kinds of agents are more markov. In comparison, we find a significant decrease in cross entropy with up to 2 seconds of context for vehicles, which is double the standard context length used for vehicles on motion prediction benchmarks (Ettinger et al., 2021; Caesar et al., 2019).
@@
figureContext Length We plot the negative log-likelihood (NLL) when we vary the context length at test-time relative to the NLL at full context. Matching with intuition, while pedestrians and cyclists are more markov on a short horizon, interaction occurs on a longer timescale for vehicles.
@@
Figure 17: Full controll rollouts Additional visualizations of full control samples from our model. The model captures the collective behavior of agents at an intersection and maneuvers such as U-turns.

### 14 Apr 2024 v2

# Trajeglish: Traffic Modeling as Next-Token Prediction

@@ Abstract
A longstanding challenge for self-driving development is simulating dynamic driving scenarios seeded from recorded driving logs. In pursuit of this functionality, we apply tools from discrete sequence modeling to model how vehicles, pedestrians and cyclists interact in driving scenarios. Using a simple data-driven tokenization scheme, we discretize trajectories to centimeter-level resolution using a small vocabulary. We then model the multi-agent sequence of discrete motion tokens with a GPT-like encoder-decoder that is autoregressive in time and takes into account intra-timestep interaction between agents. Scenarios sampled from our model exhibit state-of-the-art realism; our model tops the Waymo Sim Agents Benchmark, surpassing prior work along the realism meta metric by 3.3% and along the interaction metric by 9.9%. We ablate our modeling choices in full autonomy and partial autonomy settings, and show that the representations learned by our model can quickly be adapted to improve performance on nuScenes. We additionally evaluate the scalability of our model with respect to parameter count and dataset size, and use density estimates from our model to quantify the saliency of context length and intra-timestep interaction for the traffic modeling task.
@@ 1 Introduction
In this work, we propose an autoregressive model of the motion of road users that can be used to simulate how humans might react if a self-driving system were to choose a given sequence of actions. At test time, as visualized in Fig. 1, the model functions as a policy, outputting a categorical distribution over the set of possible states an agent might move to at each timestep. Iteratively sampling actions from the model results in diverse, scene-consistent multi-agent rollouts of arbitrary length. We call our approach Trajeglish (“tra-JEG-lish”) due to the fact that we model multi-agent trajectories as a sequence of discrete tokens, similar to the representation used in language modeling, and to make an analogy between how road users use vehicle motion to communicate and how people use verbal languages, like English, to communicate.
@@ 1.1 Related Work
Figure 2: Trajeglish Visualizations of samples from our model. Rollouts within each row are given the same single-timestep initialization, outlined in black. Future trajectories become lighter for timesteps farther into the future. Note that while some tracks overlap in the figure, they do not overlap when time is taken into account; there are no collisions in these rollouts. Videos are available on our project page.
@@ 3.1 Tokenization
We propose an easily parallelizable approach for finding template sets with low discretization error. We collect a large number of state transitions observed in data, sample one of them, filter transitions that are within $\epsilon$ meters, and repeat $|V|$ times. Pseudocode for this algorithm is included in Alg. 1. We call this method for sampling candidate templates “k-disks” given its similarity to k-means++, the standard algorithm for seeding the anchors k-means (Arthur & Vassilvitskii, 2007), as well as the Poisson disk sampling algorithm (Cook, 1986). We visualize the template sets found using k-disks with minimum discretization error in Fig. 4. We verify in Fig. 5 that the tokenized action distribution is similar on WOMD train and validation despite the fact that the templates are optimized on the training set. We show in Fig. 6 that the discretization error induced by templates sampled with k-disks is in general much better than that of k-means, across agent types. A comprehensive evaluation of k-disks in comparison to baselines is in Sec. A.3.
Figure 7: Trajeglish modeling We train an encoder-decoder transformer that predicts the action token of an agent conditional on context such as previously selected action tokens, map information, and initial agent states. The diagram visualizes the forward pass of the network during training in which initial agent states and map objects are passed into the network, and the model is trained with standard LLM-like next-token prediction on the sequence of multi-agent action tokens, shown in the top right. The bolded components are transformers.
@@ 3.2 Modeling
Our model takes as input two modalities that encode the initial scene. The first is the initial state of the agents in the scene which includes the length, width, initial position, initial heading, and object class. We apply a single-layer MLP to encode these values per-agent to an embedding of size $C$. We then add a positional embedding that encodes the agent’s order as well as agent identity across the action sequence. The second modality is the map. We use the WOMD representation of a map as a collection of “map objects”, where a map object is a variable-length polyline representing a lane, a sidewalk, or a crosswalk, for example. We apply a VectorNet encoder to encode the map to a sequence of embeddings for at most $M$ map objects (Gao et al., 2020). Note that although the model is not permutation equivariant to the agents, it is permutation invariant to the ordering of the map objects. Similar to Wayformer (Nayakanti et al., 2022), we then apply a layer of latent query attention that outputs a final encoding of the scene initialization.
@@ 4.1 WOMD Sim Agents Benchmark
Figure 8: Partial control ADE Left shows the ADE for the vehicles selected for evaluation under partial control, but for rollouts where the agents are fully autonomous. Right shows the ADE for the same vehicles but with all other agents on replay. When agents controlled by Trajeglish go first in the permutation order, they behave similarly to the no intra model. When they go last, they utilize the intra-timestep information to produce interaction more similar to recorded logs, achieving a lower ADE.
@@ 4.2 Ablation
To simplify our ablation study, we test models in this section on the scenarios they train on, of at most 24 agents and 6.4 seconds in length. We compare performance across 5 variants of our model. Both “trajeglish” and “trajeglish w/ reg.” refer to our model, the latter using the noisy tokenization strategy discussed in Sec. A.2. The “no intra” model is an important baseline designed to mimic the behavior of behavior cloning baselines used in Symphony (Igl et al., 2022) and “Imitation Is Not Enough” (Lu et al., 2023). For this baseline, we keep the same architecture but adjust the masking strategy in the decoder to not attend to actions already chosen for the current timestep. The “marginal” baseline is designed to mimic the behavior of models such as Wayformer (Nayakanti et al., 2022) and MultiPath++ (Varadarajan et al., 2021) that are trained to model the distribution over single-agent trajectories instead of multi-agent scene-consistent trajectories. For this baseline, we keep the same architecture but apply a mask to the decoder that enforces that the model can only attend to previous actions chosen by the current agent. Our final baseline is the same as the marginal baseline but without a map encoder. We use this baseline to understand the extent to which the models rely on the map for traffic modeling.
@@
We report results in Fig. 8 in a partial controllability setting in which a single agent in each scenario is chosen to be controlled by the traffic model and all other agents are set to replay. The single-agent ADE (average distance error) for the controlled-agent is similar in full autonomy rollouts for all models other than the model that does not condition on the map, as expected. However, in rollouts where all other agents are placed on replay, the replay trajectories leak information about the trajectory that the controlled-agent took in the data, and as a result, the no-intra and trajeglish rollouts have a lower ADE. Additionally, the Trajeglish rollouts in which the controlled agent is placed first do not condition on intra-timestep information and therefore behave identically to the no-intra baseline, whereas rollouts where the controlled-agent is placed last in the order provide the model with more information about the replay trajectories and result in a decreased ADE.
@@ 4.3 Analysis
We visualize the embeddings that the model learns in Fig. 11. Through the task of predicting the next token, the model learns a similarity matrix across tokens that reflects the Euclidean distance between the raw actions that the tokens represent.
@@
Figure 11: Token Embedding Visualization We run PCA on the model embeddings at initialization and at convergence, and plot the $(x,y)$ location of each of the token templates using the top 3 principal component values to determine the hue, lightness, and saturation of the point. The model learns that tokens that correspond to actions close together in euclidean space represent semantically similar actions. Note that the heading of each action is not visualized, which also affects action similarity, especially at low speeds. Additionally, the top 3 principal components include only 35% of the variance, explaining why some colors repeat.
@@ A.3 Tokenization Analysis
$(x,y,h)$-grid - We independently discretize change in longitudinal and lateral position and change in heading, and treat the template set as the product of these three sets. For vocabulary sizes of 128/256/384/512 respectively, we use 6/7/8/9 values for $x$ and $y$, and 4/6/7/8 values for $h$. These values are spaced evenly between (-0.3, 3.5) m for $x$, (-0.2 m, 0.2 m) for $y$, and (-0.1, 0.1) rad for $h$.
$(x,y)$-grid - We independently discretize change in only the location. We choose the heading for each template based on the heading of the state-to-state transition found in the data with a change in location closest to the template location. Compared to the $(x,y,h)$-grid baseline, this approach assumes heading is deterministic given location in order to gain resolution in location. We use 12/16/20/23 values for $x$ and $y$ with the same bounds as in the $(x,y,h)$-grid baseline.
@@
We offer several comparisons between these methods. In Fig. 12, we plot the expected corner distance between trajectories and tokenized trajectories as a function of trajectory length for the template sets found with k-disks. In Fig. 13, we compare the tokenization error as a function of trajectory length and find that grid-based tokenizers create large oscillations. To calibrate to a metric more relevant to the traffic modeling task, we compare the collision rate between raw trajectories as a function of trajectory length for the raw scenarios and the tokenized scenarios using k-disk template sets of size 128, 256, 384, and 512 in Fig. 14. We observe that a vocabulary size of 384 is sufficient to avoid creating extraneous collisions. Finally, Fig. 15 plots the full distribution of discretization errors for each of the baselines and Tab. 2 reports the expected discretization error across vocabulary sizes for each of the methods.
@@ A.5 Extended Rollouts for WOMD Sim Agents Benchmark
Table 3: 2023 WOMD sim agents validation
@@
A.6 December 28, 2023 - Updated Sim Agents Metrics
On December 28, 2023, Waymo announced an adjustment to the metrics for the Sim Agents benchmark to improve accuracy of vehicle and off-road collision checking (more details about this adjustment can be found here). Upon re-optimizing hyperparameters of Trajeglish for the new metrics, we found that the optimal sampling hyperparameters were $\tau=1.0$ and $p_{\mathrm{top}}=1.0$, which is more intuitive than our previously chosen hyperparameter of $\tau=1.5$ given that the metrics are intended to measure the extent to which the distribution of sampled scenarios and recorded scenarios match. We then re-trained our model to condition on 32 agents at a time instead of 24 which also improved results slightly. For the final leaderboard results before the announcement of the 2024 Sim Agents Challenge, Trajeglish did end up ahead of all models it had beaten under the previous metrics, although by much slimmer margins, shown in Tab. 4.
Table 4: WOMD sim agents validation - updated metrics
Method Realism Meta metric $\uparrow$ Kinematic metrics $\uparrow$ Interactive metrics $\uparrow$ Map-based metrics $\uparrow$ Trajeglish ($\tau=1.5$) 0.6078 0.4019 0.7274 0.7682 MTR_E 0.6348 0.4180 0.7416 0.8400 MVTA 0.6361 0.4175 0.7543 0.8253 Trajeglish ($\tau=1.0$) 0.6437 0.4157 0.7816 0.8213 MVTE 0.6448 0.4202 0.7666 0.8387 Trajeglish ($\tau=1.0$, AA=32) 0.6451 0.4166 0.7845 0.8216
A.7 Additional Ablation Results
@@ A.8 Additional Analysis
Context length refers to the number of tokens that the model has to condition on when predicting the distribution over the next token. Intuitively, as the model is given more context, the model should get strictly better at predicting the next token. We quantify this effect in Fig. A.8. We find that the relative decrease in cross entropy from increasing the context length drops off steeply for our model for pedestrians and cyclists, which aligns with the standard intuition that these kinds of agents are more Markovian. In comparison, we find a significant decrease in cross entropy with up to 2 seconds of context for vehicles, which is double the standard context length used for vehicles on motion prediction benchmarks (Ettinger et al., 2021; Caesar et al., 2019).
@@
figureContext Length We plot the negative log-likelihood (NLL) when we vary the context length at test-time relative to the NLL at full context. Matching with intuition, while pedestrians and cyclists are more Markovian on a short horizon, interaction occurs on a longer timescale for vehicles.
@@
Figure 17: Full control rollouts Additional visualizations of full control samples from our model. The model captures the collective behavior of agents at an intersection and individual maneuvers such as U-turns.